(b): Determine the mass and the weight of the al cu dimensions are 6 m 6m 8 m. Assume the density of the air is 1.16 kg/m (02) 2-2 (a) Explain why some people experience nose bleeding and some others axperience shortness of breath at high elevations (02) (b) Determine the atmospheric pressure at a location where the barometric (02) heari ior a fiuid stream reading is 750 mmHg. Take the sp gravity of mercury to be 13.6


plz answer both of these
Answer #1

a)Some people experience nose bleeding because at high altitudes atmospheric pressure is less and pressure inside body is high in comparision to atmospheric pressure so due to pressure difference it happens and air also become dry this cause drying of nasal mucosa and blood vessel get ruptured easily

Where as some people experience shortness of breath due to less oxygen

b) pressure =density of mercury ×g×h

Pressure=13.6×10^3×9.8×750×10^-3=99.96×10^3 pascal